# José Luis Blanco-Claraco

### v1

@@ 3.12 Georeferencing view-based maps
Figure 10: The problem of georeferencing a simple-map involves (a) finding the unknown relative pose between the ENU and metric map frames of reference, which can be turned up into optimizing the factor graph in (b). See discussion in §3.12.
@@
In particular, we address the problem of georeferencing view-based maps $\mathfrak{m}$ (see Eq. (24)–25), or segments of such maps, that contain frames with observations $\mathfrak{o}_{i,j}$ from low-cost GNSS receivers with positioning errors in the range of meters. Apart of its utility on itself, georeferencing segments of large simple-maps becomes a valuable tool for loop-closure detection outdoors (see §3.13).
@@
Therefore, note that once we have run LO and have GNSS observations, two coordinate origins coexist: the LO metric map origin (“map”) and that for ENU coordinates (“enu”). Georeferencing the map hence becomes finding out the rigid transformation $\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{T}}\prescript{}{map}{}\in SE(3)$ that best explains all observations, as sketched in Figure 10(a). The placement of the ENU frame may be initialized from the very first GNSS observation in the view-based map, or may be given already by another past observation in the context of sub-mapping (see LC in §3.13). We propose using factor graphs (Murphy (2012)), a popular kind of bipartite graphical model, as the ideal tool to solve such estimation problem from noisy, possibly containing spurious, GNSS readings. As shown in Figure 10(b), the graph unknowns are the relative pose $\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{T}}\prescript{}{map}{}$ and the vehicle trajectory poses $\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{T}}\prescript{}{i}{}$, in the ENU frame. The relative pose between the latter ones will normally not change significantly, since they were estimated using LO with a small uncertainty, but their values are left as free variables such as, for long trajectories, global GNSS observations are able to correct the LO drift, up to the GNSS receiver accuracy. Two types of factors are defined. First, unary GNSS factors ($f_{G}$), whose error vector $\mathbf{e}_{G}(\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{T}}\prescript{}{i}{})$ is the mismatch between the measured GNSS ENU coordinates $\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{t}}\prescript{}{i}{}$ and the predicted position of the GNSS antenna:
@@
We use GTSAM (Dellaert (2021)) to implement and solve the optimization problem defined by this factor graph. The quality of the resulting georeferencing is assessed via the square roots of the diagonal of the marginal covariance matrix for $\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{T}}\prescript{}{map}{}$, which reflect whether all six degrees of freedom in SE(3) were properly constrained by observations or not, e.g. a common caveat is the “roll” angle not be estimated accurately when processing a map segment with an almost perfect straight line trajectory. Adding multiple GNSS sensors with a large baseline could be interesting for large vehicles. Experimental results regarding georeferencing maps are shown in §5.2, while this feature is also a building block of LC, addressed next.
@@ 3.13 Loop closure algorithm
Figure 11: The two-level hierarchical graph used to determine loop closures and optimize the resulting key-frames. On the bottom, all map key-frames; on a higher level, one single frame per sub-map. Some sub-maps may include GNSS observations, allowing the determination of their coordinates in an ENU frame of reference. Refer to discussion in §3.13.
In the following, please check Algorithm 2 to follow the discussion on the top-level structure of the LC procedure. First of all, the input and output of LC are both the view-based map $\mathfrak{m}$ (see definition in §3.10), comprising key-frames with SE(3) poses with respect to some arbitrary origin and raw sensory data (i.e. GNSS observations, LiDAR scans). The goal of LC is to detect when key-frames correspond to observations of the same place, even if their global coordinates from LO are far apart, and to correct all poses so the trajectory, and consequently the associated metric maps, are globally consistent. The problem is attacked from a hierarchical view-point by first splitting the input map into sub-maps (line 3 in the pseudocode), for which their local metric map bounding box (line 5) and geo-referenced location from GNSS data (line 6, see §3.12) are evaluated. Note that GNSS is totally optional, but it will indeed help constraining the posterior search for LC candidates. Then, we build a two-level graph structure: on the bottom-level, a key-frame graph ($G_{k}$) has one node per map key-frame, while on a higher level, a sub-map graph ($G_{s}$) has one node per sub-map, as sketched in Figure 11.
@@
Using two graphs is motivated by their distinct use: the low-level graph is tightly coupled with a corresponding factor-graph in charge of actually optimizing the global poses for the entire map, while the higher-level sub-map graph is used to efficiently search for LC candidates. Initially, the low-level (key-frames) graph has a linear connection pattern, that is, each node $i$ is connected via a pose constraint ($\hskip 0.80002pt\prescript{}{i-1}{\boldsymbol{T}}\prescript{}{i}{}$) to the immediately previous node $i-1$. Exactly the same happens on the sub-map (higher) level, where we use the notation ${}_{i-1}\mathbf{P}_{i}$ to denote relative poses between sub-maps. These two graphs are built in lines 8–9 of Algorithm 2. Then, if there is more than one sub-map where georeferencing was successful, we can exploit that information by means of additional connections between the reference key-frames of the first ever sub-map with GNSS data and all others; those additional edges are labeled as ${\hskip 0.80002pt\prescript{}{b}{\boldsymbol{T}}\prescript{}{a}{}}^{GNSS}$ and ${{}_{b}\mathbf{P}_{a}}^{GNSS}$ in Figure 11 for the lower and higher levels, respectively. If there are such additional GNSS constraints, the key-frame graph is optimized using a factor graph (lines 10–12 of the pseudocode) to reduce absolute coordinate errors from those accumulated by pure LO (unbounded), to those of the GNSS receivers (maybe several meters, but bounded). It is worth mentioning that optimization is run in two passes: a first one without robust kernels, then a second robust optimization. This is to enable convergence in presence of extremely large initial global pose errors.
@@
Then, lines 13–22 comprise the loop to search for potential LC candidates, evaluate and apply them if found plausible, re-optimize all poses, and repeat until no further feasible candidates are found. A LC candidate is a pair of sub-maps that may correspond to the same physical place, at least, partially. Potential candidates are determined by evaluating, using Montecarlo to cope with uncertainty, the volumetric intersection ratio between the bounding boxes of every pair of sub-maps and accepting those above a given threshold (line 14). To find out the probabilistic relative pose between all sub-maps, we apply Dijkstra’s algorithm to the higher-level graph to obtain, for each sub-map, a tree reflecting the shortest paths to all other sub-maps. Relative poses are then computed by composing transformations following the tree edges. Note that adding GNSS edges makes that sub-maps that initially may be far from each other (large uncertainty in their relative pose) now are topologically closer, hence their relative pose based on GNSS will be used instead of the pose composition chain throughout the lower key-frame level. This graph-based procedure is similar to many past works, such as Bosse et al. (2004) or Blanco et al. (2008).
@@
Each such potential LC candidate is then actually evaluated (line 19). Our approach is based on running a configurable ICP pipeline on the metric maps built from the view-based maps of the sub-maps, including the final evaluation of the registration quality (see §3.7.4) to decide whether the LC seems plausible or not. If it does, new edges are added to both graph levels with the obtained relative pose.
@@
A few words are in order regarding this ICP pipeline, which is different than that in Figure 6 for LO. The first difference is the use of two point-cloud map layers: one for all objects near the ground plane, and another for everything else. The ICP pipeline then includes two instances of point-to-point matchers such that “ground” and “non-ground” points are only attempted to be paired against points belonging to the same category in the other sub-maps. On ICP quality assessment, here we include the voxel occupancy-based method described in §3.7.4. Furthermore, the point-cloud metric maps used for LC have a different implementation than that used during LO: instead of the hashed-voxels maps used for LO, we use simple clouds with contiguous memory layout. The main motivation for this is that search for nearest neighbors (NN) (one of the dominating costs in ICP algorithms) with large search radii is more efficient using KD-Trees (the implementation of NN for the latter map type) than searching neighboring voxels in hashed maps.
@@
Newer College Dataset $\rightarrow$ § 4.9
@@
OS0-128 128 Handheld College 12 78.2k 9.8 km
Newer College Dataset $\rightarrow$ § 4.9
@@
OS1-64 Handheld College 2 41.9k 4.7 km
MulRan $\rightarrow$ § 4.1
@@
OS1-64 64 Vehicle Campus/City 12 127.7k 123.4 km
KITTI $\rightarrow$ § 4.2
@@
HDL-64E Vehicle Road/Urban/Mixed 22 43.5k 44.24 km
Voxgraph $\rightarrow$ § 4.7
@@
OS1-64 Drone Urban 1 2.2k 0.24 km
HILTI 2021 $\rightarrow$ § 4.6
@@
OS0-64 Drone Industrial unit 1 0.9k 0.08 km
KITTI-360 $\rightarrow$ § 4.3
@@
HDL-64E Vehicle Urban 12 70.1k 58.7 km
ParisLuco $\rightarrow$ § 4.4
@@
HDL32 Vehicle City 1 12.7k 4.0 km
Almería forests $\rightarrow$ § 5.3
@@
OS0-32 32 Backpack Forests 6 19.9k 1.6 km
NTU-VIRAL $\rightarrow$ § 4.5
@@
2 $\times$ OS1-16 Drone Campus 9 32.6k 2.2 km
DARPA Subterranean $\rightarrow$ § 4.8
@@
VLP16 16 ANYmal C Caves 4 143.1k 2.0 km
UAL campus $\rightarrow$ § 4.10
@@
SICK LRF 1 PowerBot Offices 1 1.2k 0.1 km
Málaga CS faculty $\rightarrow$ § 5.1
@@
SICK LRF Wheelchair Campus 1 4.7k 1.9 km
@@ 4 LiDAR odometry quantitative experimental validation
Please, refer to Table 1 for an overview of all the used datasets. Note that we selected an extensive list of public datasets in order to cover a large variety of motion models (vehicles, drones, hand-held, etc.) and sensor models and resolutions. In particular, it is noteworthy that we used exactly the same system configuration for all 3D LiDAR datasets, from 16 to 128 beams. In all cases, our system provided accurate trajectory estimations without diverging, while keeping execution times fast enough for running significantly faster than the real sensor rate.
@@
  • Absolute Trajectory Error (ATE) (m): We use the RMSE value of the distances between time-aligned ground-truth and estimated trajectories, after spatially aligning them using Umeyama’a method (Umeyama (1991)). Evaluation is performed using the open-sourced tool evo_traj (Grupp (2017)).
Unless noted otherwise, metrics for alternative methods have been taken from their original publications. The exception is KISS-ICP (Vizzo et al. (2023)), which due to its simplicity of integration into third-party systems, has been integrated as another MOLA LO module, hence enabling running all datasets on it and obtaining trajectories in the same format than our proposed system, becoming a excellent baseline SOTA reference for all the datasets.
@@
Figure 12: Estimated trajectories for the proposed LiDAR odometry system applied to the MulRan dataset, compared to ground truth. Only one sequence is shown for each of the four locations. Estimated paths including loop closure (section 3.13) are denoted with legend “MOLA-LO+LC”.

### v2

@@ 3.11 Georeferencing view-based maps
Figure 11: The problem of georeferencing a simple-map involves (a) finding the unknown relative pose between the ENU and metric map frames of reference, which can be turned up into optimizing the factor graph in (b). See discussion in Section 3.11.
@@
In particular, we address the problem of georeferencing view-based maps $\mathfrak{m}$ (see Eq. (24)–25), or segments of such maps, that contain frames with observations $\mathfrak{o}_{i,j}$ from low-cost GNSS receivers with positioning errors in the range of meters. Apart of its utility on itself, georeferencing segments of large maps becomes a valuable tool for loop-closure detection outdoors (see Section 3.12).
@@
Therefore, note that once we have run LO and have GNSS observations, two coordinate origins coexist: the LO metric map origin (“map”) and that for ENU coordinates (“enu”). Georeferencing the map hence becomes finding out the rigid transformation $\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{T}}\prescript{}{map}{}\in SE(3)$ that best explains all observations, as sketched in Figure 11(a). The placement of the ENU frame may be initialized from the very first GNSS observation in the view-based map, or may be given already by another past observation in the context of sub-mapping (see LC in Section 3.12). We propose using factor graphs (Ni et al., 2007; Murphy, 2012), a popular kind of bipartite graphical model, as the ideal tool to solve such estimation problem from noisy, possibly containing spurious, GNSS readings. As shown in Figure 11(b), the graph unknowns are the relative pose $\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{T}}\prescript{}{map}{}$ and the vehicle trajectory poses $\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{T}}\prescript{}{i}{}$, in the ENU frame. The relative pose between the latter ones will normally not change significantly, since they were estimated using LO with a small uncertainty, but their values are left as free variables such as, for long trajectories, global GNSS observations are able to correct the LO drift, up to the GNSS receiver accuracy. In particular, GNSS observations are specially beneficial to reduce the larger error that exists in the vertical axis due to LO drift in long trajectories. Two types of factors are defined. First, unary GNSS factors ($f_{G}$), whose error vector $\mathbf{e}_{G}(\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{T}}\prescript{}{i}{})$ is the mismatch between the measured GNSS ENU coordinates $\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{t}}\prescript{}{i}{}$ and the predicted position of the GNSS antenna:
@@
We use GTSAM (Dellaert (2021)) to implement and solve the optimization problem defined by this factor graph. The quality of the resulting georeferencing is assessed via the square roots of the diagonal of the marginal covariance matrix for $\hskip 0.80002pt\prescript{}{enu}{\boldsymbol{T}}\prescript{}{map}{}$, which reflect whether all six degrees of freedom in SE(3) were properly constrained by observations or not, e.g. a common caveat is the “roll” angle not be estimated accurately when processing a map segment with an almost perfect straight line trajectory. Adding multiple GNSS sensors with a large baseline could be interesting for large vehicles. Experimental results regarding georeferencing maps are shown in Section 5.2, while this feature is also a building block of LC, addressed next.
@@ 3.12 Loop closure algorithm
Figure 12: The two-level hierarchical graph used to determine loop closures and optimize the resulting key-frames. On the bottom, all map key-frames; on a higher level, one single frame per sub-map. Some sub-maps may include GNSS observations, allowing the determination of their coordinates in an ENU frame of reference. Refer to discussion in Section 3.12.
@@
In the following, please check Algorithm 2 to follow the discussion on the top-level structure of the LC procedure. First of all, the input and output of LC are both the view-based map $\mathfrak{m}$ (see definition in Section 3.9), comprising key-frames with SE(3) poses with respect to some arbitrary origin and raw sensory data (i.e. GNSS observations, LiDAR scans). The goal of LC is to detect when key-frames correspond to observations of the same place, even if their global coordinates from LO are far apart, and to correct all poses so the trajectory, and consequently the associated metric maps, are globally consistent. The problem is attacked from a hierarchical view-point by first splitting the input map into sub-maps (line 3 in the pseudocode), for which their local metric map bounding box (line 5) and geo-referenced location from GNSS data (line 6, see Section 3.11) are evaluated. Note that GNSS is totally optional, but it will indeed help constraining the posterior search for LC candidates by reducing the uncertainty of relative poses between key-frames that are topologically far apart. Then, we build a two-level graph structure: on the bottom-level, a key-frame graph ($G_{k}$) has one node per map key-frame, while on a higher level, a sub-map graph ($G_{s}$) has one node per sub-map, as sketched in Figure 12. This approach follows a tradition of splitting large metric maps into submaps, such as done with Tectonic SAM in Ni et al. (2007) or with condensed factors in Grisetti et al. (2012). Two relevant differences of our approach with respect to those past works are: (i) the explicit distinction between condensed factors coming from LiDAR odometry and those from GNSS observations, and (ii) the usage of each of the two graphs for different purposes. In particular, the low-level graph is tightly coupled with a corresponding factor-graph in charge of actually optimizing the global poses for the entire map, while the higher-level sub-map graph is used to efficiently search for LC candidates. Initially, the low-level (key-frames) graph has a linear connection pattern, that is, each node $i$ is connected via a pose constraint ($\hskip 0.80002pt\prescript{}{i-1}{\boldsymbol{T}}\prescript{}{i}{}$) to the immediately previous node $i-1$. Exactly the same happens on the sub-map (higher) level, where we use the notation ${}_{i-1}\mathbf{P}_{i}$ to denote relative poses between sub-maps. These two graphs are built in lines 8–9 of Algorithm 2. Then, if there is more than one sub-map where georeferencing was successful, we can exploit that information by means of additional connections between the reference key-frames of the first ever sub-map with GNSS data and all others; those additional edges are labeled as ${\hskip 0.80002pt\prescript{}{b}{\boldsymbol{T}}\prescript{}{a}{}}^{GNSS}$ and ${{}_{b}\mathbf{P}_{a}}^{GNSS}$ in Figure 12 for the lower and higher levels, respectively. If there are such additional GNSS constraints, the key-frame graph is optimized using a factor graph (lines 10–12 of the pseudocode) to reduce absolute coordinate errors from those accumulated by pure LO (unbounded), to those of the GNSS receivers (maybe several meters, but bounded). It is worth mentioning that optimization is run in two passes: a first one without robust kernels, then a second robust optimization. Note that this is the opposite of what is typically done with SLAM frontends, where robust optimization comes first to reject outliers. However, if we want to handle loop closures with a large initial mismatch between the global poses of the two involved submaps (e.g. thanks to GNSS data, if available) robust kernels cannot be applied at first, or the error would be so large that it will be ignored as if it was an outlier.
@@
Then, lines 13–22 comprise the loop to search for potential LC candidates, evaluate and apply them if found plausible, re-optimize all poses, and repeat until no further feasible candidates are found. A LC candidate is a pair of sub-maps that may correspond to the same physical place, at least, partially. Potential candidates are determined by evaluating the volumetric intersection ratio between the bounding boxes of every pair of sub-maps and accepting those above a given threshold (line 14). Given the potential large uncertainty in the relative poses of two submaps, metrically close to each other after a long topological loop, we use a Monte Carlo method to draw relative pose samples to evaluate the mathematical expectation of the intersection ratio. To find out the probabilistic relative pose between all sub-maps, we apply Dijkstra’s algorithm to the higher-level graph to obtain, for each sub-map, a tree reflecting the shortest paths to all other sub-maps. Relative poses are then computed by composing transformations following the tree edges. Note that adding GNSS edges makes that sub-maps that initially may be far from each other (large uncertainty in their relative pose) now are topologically closer, hence their relative pose based on GNSS will be used instead of the pose composition chain throughout the lower key-frame level. This graph-based procedure is similar to many past works, such as Bosse et al. (2004) or Blanco et al. (2008). Note that an alternative way to determine potential loop closure candidates would be using descriptors for local maps, as classically done in visual place recognition or, in the context of LiDAR SLAM, in methods such as Scan Context (Kim and Kim, 2018) or ScanContext++ (Kim et al., 2021).
@@
Each such potential LC candidate is then actually evaluated (line 19). Our approach is based on running a configurable ICP pipeline on the metric maps built from the view-based maps of the sub-maps, including the final evaluation of the registration quality (see Section 3.6.4) to decide whether the LC seems plausible or not. If it does, new edges are added to both graph levels with the obtained relative pose.
A few words are in order regarding this ICP pipeline, which is different than that in Figure 6 for LO. The first difference is the use of two point-cloud map layers: one for all objects near the ground plane, and another for everything else. The ICP pipeline then includes two instances of point-to-point matchers such that “ground” and “non-ground” points are only attempted to be paired against points belonging to the same category in the other sub-maps. Even a naive method to classify the point cloud into these two categories such as split by $z$ (vertical) coordinates leads to good results, as long as local maps do not have significant tilt and the ground floor can be successfully segmented. More advanced methods have been used in the literature to segment the ground points in the context of loop closure, such as in Quatro++ (Lim et al., 2024), and could be used instead of our naive classifier.
@@
Regarding ICP quality assessment, here we include the voxel occupancy-based method described in Section 3.6.4. Furthermore, the point-cloud metric maps used for LC have a different implementation than that used during LO: instead of the hashed-voxels maps used for LO, we use simple clouds with contiguous memory layout. The main motivation for this is that search for nearest neighbors (NN) (one of the dominating costs in ICP algorithms) with large search radii is more efficient using KD-Trees (the implementation of NN for the latter map type) than searching neighboring voxels in hashed maps.
@@
Newer College Dataset $\rightarrow$ Section 4.9
@@
OS0-128 128 - Handheld College 12 78.2k 9.8 km
Newer College Dataset $\rightarrow$ Section 4.9
@@
OS1-64 - Handheld College 2 41.9k 4.7 km
MulRan $\rightarrow$ Section 4.1
@@
OS1-64 Vehicle Campus/City 12 127.7k 123.4 km
KITTI $\rightarrow$ Section 4.2
@@
HDL-64E 64 - Vehicle Road/Urban 22 43.5k 44.24 km
Voxgraph $\rightarrow$ Section 4.7
@@
OS1-64 - Drone Urban 1 2.2k 0.24 km
HILTI 2021 $\rightarrow$ Section 4.6
@@
OS0-64 - Drone
Industrial
unit
1 0.9k 0.08 km
KITTI-360 $\rightarrow$ Section 4.3
@@
HDL-64E - Vehicle Urban 12 70.1k 58.7 km
ParisLuco $\rightarrow$ Section 4.4
@@
HDL32 32 - Vehicle City 1 12.7k 4.0 km
Almería forests $\rightarrow$ Section 5.3
@@
OS0-32 - Backpack Forests 6 19.9k 1.6 km
NTU-VIRAL $\rightarrow$ Section 4.5
@@
2 $\times$ OS1-16 - Drone Campus 9 32.6k 2.2 km
DARPA Subterranean $\rightarrow$ Section 4.8
@@
VLP16 16 - ANYmal C Caves 4 143.1k 2.0 km
UAL campus $\rightarrow$ Section 4.10
@@
SICK LRF 1 - PowerBot Offices 1 1.2k 0.1 km
Málaga CS faculty $\rightarrow$ Section 5.1
@@
SICK LRF - Wheelchair Campus 1 4.7k 1.9 km
@@
3.13 Dataset sources
Finally, it is worth mentioning that the former work in Blanco (2019) already introduced data source MOLA modules as an abstraction of either, real live sensors or offline datasets. Such abstraction aims at enabling algorithm implementations in MOLA to immediately work with any of the supported datasets, without worrying about the details of each dataset file and data structures, sensor intrinsic and extrinsic calibration, ground truth (GT) axes conventions, etc.
  • MRPT rawlog files: Rawlog files are an equivalent to ROS bag files, introduced in the MRPT framework in 2005. It has been used in public datasets, e.g. Blanco-Claraco et al. (2014, 2019). Backwards compatibility and Operative System independence has been ensured during the whole life of the project, hence software written today using the latest MRPT version is able to read all past datasets, even if generated in a different OS or processor architecture. There exists an import tool from ROS 1 bag files to rawlogs, hence this MOLA input module also offers easy access to many of the ROS 1 public datasets, e.g. HILTI challenge datasets (Helmberger et al., 2022) or the Newer College dataset (Zhang et al., 2021).
Finally, a word is in order regarding memory management for dataset input modules. Loading and keeping in memory whole datasets may become unfeasible for a typical desktop computer, e.g. some individual sequences of the MulRan dataset (Kim et al., 2020) take 29 GiB. The present framework exploits the usage of MRPT lazy-load mechanism in different areas, including dataset sources. Observations that tend to be heavy in memory requirements (i.e. images, 3D LiDAR scans, and RGB+D frames) can be automatically loaded to memory when accessed, to be unloaded once a fixed number of additional frames have been processed. Most MOLA dataset input modules also feature read-ahead in a multi-thread fashion to minimize the performance impact of such lazy-load mechanism. In practice, this creates a cache of recently-used and about-to-be-used sensory data quickly accessible in RAM while SLAM is processing them.
@@ 4 LiDAR odometry quantitative experimental validation
Please, refer to Table 1 for an overview of all the used datasets. Note that we selected an extensive list of public datasets in order to cover a large variety of motion models (vehicles, drones, hand-held, etc.) and sensor models and resolutions. In particular, it is noteworthy that we used exactly the same system configuration (called “default configuration” in the following) for all 3D LiDAR datasets, from 16 to 128 beams. In all cases, our system provided accurate trajectory estimations without diverging, while keeping execution times fast enough for running significantly faster than the real sensor rate. That said, these additional configurations have been also evaluated for selected datasets with the aim of serving for ablation studies in Section 7, to demonstrate a particular feature, or to illustrate the flexibility of the proposed system:
  • A 2D-LiDAR configuration, addressed in Section 5.1.
  • A configuration where the local map update decider (see Section 3.4) is not used and the map is updated for all time steps.
  • A configuration using an alternative 3D-NDT metric map instead of plain point clouds.
@@
  • Absolute Trajectory Error (ATE) (m): We use the RMSE value of the distances between time-aligned ground-truth and estimated trajectories, after spatially aligning them using Umeyama’a method (Umeyama, 1991). Evaluation is performed using the open-sourced tool evo_traj (Grupp, 2017).
Unless noted otherwise, metrics for alternative methods have been taken from their original publications. The exception is KISS-ICP (Vizzo et al., 2023), which due to its simplicity of integration into third-party systems, has been integrated as another MOLA LO module, hence enabling running all datasets on it and obtaining trajectories in the same format than our proposed system, becoming a excellent baseline SOTA reference for all the datasets.
@@
Figure 13: Estimated trajectories for the proposed LiDAR odometry system (in its default configuration) applied to the MulRan dataset, compared to ground truth. Only one sequence is shown for each of the four locations. Estimated paths including loop closure (section 3.12) are denoted with legend “MOLA-LO+LC”.